OpenStudy (anonymous):

HELP....write an equation of the line containing the given point and perpendicular to the given line. (0,8); 6x+7y=2

ganeshie8 (ganeshie8):

start by finding the slope of given line : 6x+7y=2

ganeshie8 (ganeshie8):

knw how to find the slope from equation ?

OpenStudy (anonymous):

Solve for y

ganeshie8 (ganeshie8):

congrats @iambatman on turning green :)

OpenStudy (anonymous):

Thank you ^.^

OpenStudy (anonymous):

to find a slope from an equation.[ax+by+c=0] \[slope=-a \div b\]then using the formula, \[y-y _{1}= m(x-x _{1})\]u can find the equation with the provided point and the found out slope..